# On cutoff via rigidity for high dimensional curved diffusions

### v3

@@ 3.1. Preliminaries on rigidity
First, note that $\sum_{i=1}^{d}m^{(d)}_{i}=\int\sum_{i=1}^{d}x_{i}\mathrm{d}\mu^{(d)}(x)=0$ since the image law of $\mu^{(d)}$ by $(x_{1},\ldots,x_{d})\mapsto x_{1}+\cdots+x_{d}$ is a centered Gaussian.
Lower bound when $t_{d}=(1-\epsilon)t_{*}$. As per Theorem 1.5, $x_{1}+\cdots+x_{d}$ is an eigenfunction. Since its squared $L^{2}(\mu)$ norm is $d\rho^{-1}$, $f_{d}(x)=\sqrt{\rho/d}(x_{1}+\cdots+x_{d})$ is a normalized eigenfunction. Now, for both $S^{(d)}=B(m^{(d)},c\sqrt{d})$ and $S^{(d)}=m^{(d)}+[-c,c]^{d}$ we have
$\sup_{x\in S^{(d)}}|f_{d}(x)|=c\sqrt{d}.$ (3.35)
@@
$\mathrm{W}_{2}(X^{(d)}_{t_{d}},\mu^{(d)})\geq\mathrm{e}^{-\rho t_{d}}c\sqrt{d}=cd^{\epsilon/2}\xrightarrow[d\to\infty$$\infty$. (3.36)
@@
$\mathrm{W}_{2}(X_{t_{d}}^{(d)},\mu^{(d)})\leq\mathrm{e}^{-\lambda_{1}^{(d)}t_{d}}\Bigr(\sup_{x_{0}\in S^{(d)}}|x_{0}|^{2}+d/\rho\Bigr)^{1/2}\leq(c^{2}+\rho^{-1})^{1/2}d^{-\epsilon/2}\xrightarrow[d\to\infty]{}0.$ (3.38)
It remains to combine Corollary 1.3 and Corollary 1.4 to conclude the proof.
@@
The spectral gap $\lambda_{1}^{(d)}$ of $X^{(d)}$ satisfies $\lambda_{1}^{(d)}\geq\kappa^{(d)}$, thus
$\lim_{d\to\infty}\lambda_{1}^{(d)}t_{0}^{(d)}=+\infty,$ (3.39)
@@
$\liminf_{d\to\infty}\sup_{x^{(d)}_{0}\in S^{(d)}}\mathrm{H}(X^{(d)}_{(1-\varepsilon/2)t_{0}^{(d)}}\mid\mu^{(d)})\geq\frac{\eta^{2}}{2}.$ (3.40)
@@
$\mathrm{H}(X^{(d)}_{t^{\prime}}\mid\mu^{(d)})\leq\mathrm{e}^{-2\kappa^{(d)}(t^{\prime}-t)}\mathrm{H}(X^{(d)}_{t}\mid\mu^{(d)}).$ (3.41)
Using $t^{\prime}=(1-\frac{\varepsilon}{2})t_{0}^{(d)}$, $t=(1-\varepsilon)t_{0}^{(d)}$, and $\lim_{d\to\infty}\kappa^{(d)}t_{0}^{(d)}=+\infty$, we get
$\liminf_{d\to\infty}\mathrm{H}(X^{(d)}_{(1-\varepsilon)t_{0}^{(d)}}\mid\mu^{(d)})\geq\mathrm{e}^{\varepsilon\limsup_{d\to\infty}\kappa^{(d)}t_{0}^{(d)}}\frac{\eta^{2}}{2}=+\infty.$ (3.42)
For the upper bound, a careful reading of the proof of (Salez, 2025, Th. 1) shows that
$\limsup_{d\to\infty}\mathrm{H}\bigr(X^{(d)}_{(1+\frac{\varepsilon}{2})t_{0}^{(d)}}\mid\mu^{(d)}\bigr)\leq C_{\varepsilon}<\infty.$ (3.43)
Using the exponential decay of the relative entropy and $\lim_{d\to\infty}\kappa^{(d)}t_{0}^{(d)}=+\infty$, we get
$\limsup_{d\to\infty}\mathrm{H}(X^{(d)}_{(1+\varepsilon)t_{0}^{(d)}}\mid\mu^{(d)})\leq\mathrm{e}^{-\varepsilon\liminf_{d\to\infty}\kappa^{(d)}t_{0}^{(d)}}C_{\varepsilon}=0.$ (3.44)
For Wasserstein distance, the upper bound comes from the one for relative entropy via the Talagrand inequality $\kappa^{(d)}\mathrm{W}_{2}(X,\mu^{(d)})^{2}\leq 2\mathrm{H}(X\mid\mu^{(d)})$, while the lower bound comes from the regularization (3.18) with $t=\varepsilon t^{(d)}_{0}$ and the semigroup property.
Finally, for Fisher information, the lower bound comes from one for the relative entropy via the logarithmic Sobolev inequality $2\kappa^{(d)}\mathrm{H}(X\mid\mu^{(d)})\leq\mathrm{I}(X\mid\mu^{(d)})$, while the upper bound comes from the regularization (3.31) and the semigroup property.
@@
$(\mathbb{R}^{k},\left|\cdot\right|_{2},\gamma_{k,\rho^{-1}})\times(M^{\prime},g^{\prime},\mu^{\prime})=(\mathbb{R}^{k}\times M^{\prime},\left|\cdot\right|_{2}\oplus g^{\prime},\gamma_{k,\rho^{-1}}\otimes\mu^{\prime})$ (3.46)
@@
  • $k$ is the dimension of the eigenspace of -$\mathcal{L}$ associated with the eigenvalue $\lambda_{1}$
  • $\gamma_{k,\rho^{-1}}$ is the centered Gaussian law on $\mathbb{R}^{k}$ with covariance matrix $\rho^{-1}\mathrm{Id}_{k}$
@@
Under the setting of Theorem 3.5, and through the isometry that it provides, any element of $E_{1}$ is of the form $a\cdot p(x)$ with $p$ the projection on the Euclidean factor of dimension $k$, $a\in\mathbb{R}^{k}$, $|a|=\sqrt{\rho}\|f\|_{L^{2}(\mu)}$. Moreover, if $f_{i}=a_{i}\cdot p(x)$, for $i=1,2$, are two orthogonal eigenfunctions, then $a_{1}\perp a_{2}$.
@@
What follows is a broad sketch, were we focus on justifying the form of the eigenfunctions, but do not discuss in too much detail the splitting of the space, that was established in Cheng and Zhou (2017). See also Gigli et al. (2020) for a full proof in the non-smooth setting of RCD spaces.
From the integrated Bochner formula, for any $g$ in the domain of $\mathcal{L}$ we have
@@
which forces $\mathrm{Hess}(f)=0$ almost everywhere, and thus everywhere since eigenfunctions are smooth. Therefore $f$ is affine, and non-constant. In particular, $\nabla f$ is a non-trivial parallel vector field, which forces the splitting of the manifold in a product form $\mathbb{R}\times M^{\prime}$, along which $f$ is only an affine function of the first coordinate. Repeating this strategy for successive orthogonal eigenfunctions gives a splitting with a $k$-dimensional Euclidean factor. See Cheng and Zhou (2017) for full details. And once the splitting is established, we can view the eigenfunctions as functions on $\mathbb{R}^{k}$.

### v4

@@ 3.1. Preliminaries on rigidity
We drop the $(d)$ superscript. We have $\sum_{i=1}^{d}m_{i}=\int\sum_{i=1}^{d}x_{i}\mathrm{d}\mu(x)=0$ since the image of $\mu$ by $(x_{1},\ldots,x_{d})\mapsto x_{1}+\cdots+x_{d}$ is a centered Gaussian.
Lower bound when $t_{d}=(1-\epsilon)t_{*}$. As per Theorem 1.5, $x_{1}+\cdots+x_{d}$ is an eigenfunction. Since its squared $L^{2}(\mu)$ norm is $d\rho^{-1}$, $f_{d}(x)=\sqrt{\rho/d}(x_{1}+\cdots+x_{d})$ is a normalized eigenfunction. Now, for both $S=B(m,c\sqrt{d})$ and $S=m+[-c,c]^{d}$ we have
$\sup_{x\in S}|f_{d}(x)|=c\sqrt{\rho}\sqrt{d}.$ (3.35)
@@
$\sup_{x_{0}\in S}\mathrm{W}_{2}(X_{t_{d}},\mu)\geq\mathrm{e}^{-\rho t_{d}}c\sqrt{d}=cd^{\epsilon/2}\xrightarrow[d\to\infty$$\infty$. (3.36)
@@
$\sup_{x_{0}\in S}\mathrm{W}_{2}(X_{t_{d}},\mu)\leq\mathrm{e}^{-\lambda_{1}t_{d}}\Bigl(\sup_{x_{0}\in S}|x_{0}-m|^{2}+d/\rho\Bigr)^{1/2}\leq(c^{2}+\rho^{-1})^{1/2}d^{-\epsilon/2}\xrightarrow[d\to\infty]{}0.$ (3.38)
It remains to proceed as for Corollary 1.3 and Corollary 1.4 to conclude the proof.
@@
Here again, we drop the $(d)$ superscript to simplify the notation. The spectral gap $\lambda_{1}$ of $X$ satisfies $\lambda_{1}\geq\kappa$, thus
$\lim_{d\to\infty}\lambda_{1}t_{0}=+\infty,$ (3.39)
@@
$\liminf_{d\to\infty}\sup_{x_{0}\in S}\mathrm{H}(X_{(1-\varepsilon/2)t_{0}}\mid\mu)\geq 2\eta^{2}.$ (3.40)
@@
$\mathrm{H}(X_{t^{\prime}}\mid\mu)\leq\mathrm{e}^{-2\kappa(t^{\prime}-t)}\mathrm{H}(X_{t}\mid\mu).$ (3.41)
Using $t^{\prime}=(1-\frac{\varepsilon}{2})t_{0}$, $t=(1-\varepsilon)t_{0}$, and $\lim_{d\to\infty}\kappa t_{0}=+\infty$, we get
$\liminf_{d\to\infty}\sup_{x_{0}\in S}\mathrm{H}(X_{(1-\varepsilon)t_{0}}\mid\mu)\geq\mathrm{e}^{\varepsilon\limsup_{d\to\infty}\kappa t_{0}}2\eta^{2}=+\infty.$ (3.42)
For the upper bound, from (Salez, 2025, Proof of Th. 1), we get
$\limsup_{d\to\infty}\sup_{x_{0}\in S}\mathrm{H}\bigl(X_{(1+\frac{\varepsilon}{2})t_{0}}\mid\mu\bigr)\leq C_{\eta,\varepsilon}<\infty.$ (3.43)
Using the exponential decay of the relative entropy and $\lim_{d\to\infty}\kappa t_{0}=+\infty$, we get
$\limsup_{d\to\infty}\sup_{x_{0}\in S}\mathrm{H}(X_{(1+\varepsilon)t_{0}}\mid\mu)\leq\mathrm{e}^{-\varepsilon\liminf_{d\to\infty}\kappa t_{0}}C_{\eta,\varepsilon}=0.$ (3.44)
[…]
$\mathrm{H}(X_{(1-\frac{\varepsilon}{2})t_{0}}\mid\mu)\leq\frac{1}{2s}\mathrm{W}_{2}(X_{(1-\varepsilon)t_{0}},\mu)^{2},$ (3.45)
and then the lower bound for $\mathrm{H}$ and the assumption $t_{0}\to\infty$:
$\sup_{x_{0}\in S}\mathrm{W}_{2}(X_{(1-\varepsilon)t_{0}},\mu)^{2}\geq 2s\times 2\eta^{2}=2\varepsilon\eta^{2}t_{0}\xrightarrow[d\to\infty$$\infty$. (3.46)
[…]
$\mathrm{I}(X_{t_{1}}\mid\mu)\leq\frac{\mathrm{H}(X_{t_{0}^{\prime}}\mid\mu)}{t_{1}-t_{0}^{\prime}}$ (3.47)
where $t_{0}^{\prime}=(1+\frac{\varepsilon}{2})t_{0}$ and $t_{1}=(1+\varepsilon)t_{0}$, then the upper bound on $\mathrm{H}$ and the assumption $t_{0}\to\infty$:
$\sup_{x_{0}\in S}\mathrm{I}(X_{(1+\varepsilon)t_{0}}\mid\mu)\leq\frac{\sup_{x_{0}\in S}\mathrm{H}(X_{(1+\frac{\varepsilon}{2})t_{0}}\mid\mu)}{\frac{\varepsilon}{2}t_{0}}\xrightarrow[d\to\infty]{}0.$ (3.48)
@@
$(\mathbb{R}^{k_{1}},\left|\cdot\right|_{2},\gamma_{k_{1},\rho^{-1}})\times(M^{\prime},g^{\prime},\mu^{\prime})=(\mathbb{R}^{k_{1}}\times M^{\prime},\left|\cdot\right|_{2}\oplus g^{\prime},\gamma_{k_{1},\rho^{-1}}\otimes\mu^{\prime})$ (3.50)
@@
  • $k_{1}$ is the dimension of the eigenspace of -$\mathcal{L}$ associated with the eigenvalue $\lambda_{1}$
  • $\gamma_{k_{1},\rho^{-1}}$ is the centered Gaussian law on $\mathbb{R}^{k_{1}}$ with covariance matrix $\rho^{-1}\mathrm{Id}_{k_{1}}$
@@
Under the setting of Theorem 3.6, and through the isometry that it provides, any element of $E_{1}$ is of the form $a\cdot p(x)$ with $p$ the projection on the Euclidean factor of dimension $k_{1}$, $a\in\mathbb{R}^{k_{1}}$, $|a|=\sqrt{\rho}\|f\|_{L^{2}(\mu)}$. Moreover, if $f_{i}=a_{i}\cdot p(x)$, for $i=1,2$, are two orthogonal eigenfunctions, then $a_{1}\perp a_{2}$.
@@
What follows is a broad sketch, where we focus on justifying the form of the eigenfunctions, but do not discuss in too much detail the splitting of the space, that was established in Cheng and Zhou (2017). See also Gigli et al. (2020) for a full proof in the non-smooth setting of RCD spaces. From the integrated Bochner formula, for any $g$ in the domain of $\mathcal{L}$ we have
@@
which forces $\mathrm{Hess}(f)=0$ almost everywhere, and thus everywhere since eigenfunctions are smooth. Therefore $f$ is affine, and non-constant. In particular, $\nabla f$ is a non-trivial parallel vector field, which forces the splitting of the manifold in a product form $\mathbb{R}\times M^{\prime}$, along which $f$ is only an affine function of the first coordinate. Repeating this strategy for successive orthogonal eigenfunctions gives a splitting with a $k_{1}$-dimensional Euclidean factor. See Cheng and Zhou (2017) for full details. And once the splitting is established, we can view the eigenfunctions as functions on $\mathbb{R}^{k_{1}}$.